# Möbius-topological auxiliary function for $f$ electrons

### v2

@@ Abstract
where $Y_{l}^{m}(\theta,\varphi)$ are spherical harmonics constructed from the associated Legendre polynomials $P_{l}^{m}(\cos\theta)$. Here, $l$ and $m$ denote the degree and order of $P_{l}^{m}$, corresponding to the orbital angular momentum and its $z$-component. Throughout this paper, unless otherwise stated, $l\in\mathbb{N}$ and $m\in\mathbb{Z}$ satisfy $l\leq m\leq l$.
@@
Figure 2: (a) The factor $e^{i\varphi/2}$ is plotted along a circular arc of radius 3, with its vector trajectory forming a Möbius strip. (b) Auxiliary densities distribution of $f$-electron states in Nd3+ ($J=9/2$). Kets $|M\rangle$ denote the quantum states. The auxiliary densities exhibit O(2) symmetry, and the surface encloses 50% of the total density.
Symmetry analysis.— Auxiliary functions can be employed to determine energy eigenstates, provided their densities conform to the crystal-field symmetry. The eigenstate structure is dictated by rotational symmetry together with a horizontal mirror plane, while vertical mirror planes constrain the phases of the complex coefficients forming the eigenstate components.
The auxiliary density of a pure eigenstate with a single $M$ exhibits continuous O(2)(2) rotational symmetry. A superposition, however, generally breaks this symmetry, reducing it to discrete rotational symmetries. For instance, for a two-state superposition $\left|\psi\right\rangle=C_{1}\left|M_{1}\right\rangle+C_{2}\left|M_{2}\right\rangle$, the auxiliary density is:
@@
where $A$ and $B$ are independent of $\varphi$. The auxiliary density thus has $|M_{1}-M_{2}|$-fold rotational symmetry. For example, $|M_{1}-M_{2}|=3$ gives threefold symmetry; =2=2, twofold; =1=1, no rotational symmetry.
@@
Therefore, the auxiliary density exhibits $\kappa$-fold rotational symmetry if and only if the following condition hold:
$\kappa=\gcd(M_{j}-M_{k}),$ (6)
where $\gcd$ (greatest common divisor) is taken over all pairs $(j,k)$. From Eq. 6, the eigenstate whose auxiliary density $|\psi|^{2}$ exhibits $\kappa$-fold rotational symmetry can be written as
$|\psi\rangle=\sum_{j=0}^{\lfloor(2J-q)/\kappa\rfloor}C_{j}\,|J-\kappa j-q\rangle,\quad q\in\{0,1,\dots,\kappa-1\}.$ (7)
The auxiliary density is always inversion-symmetric (see proof in SM). When combined with even-fold rotational symmetry, it automatically exhibits horizontal mirror symmetry. For odd-fold rotational symmetry with inherent inversion symmetry, an introduction of a horizontal mirror plane will double the rotational order, effectively making it even-fold. In this sense, horizontal mirror symmetry amplifies odd-fold rotations while leaving even-fold rotations unchanged. Depending on $\kappa$, this enhancement changes the symmetry from Cκ to either C2κm or Cκm as:
$\mathrm{I}+\mathrm{C}_{\kappa}=\begin{cases}\mathrm{S}_{2\kappa}\xrightarrow{+\sigma_{z}}\mathrm{C}_{2\kappa\mathrm{m}},&\text{if }\kappa\text{ is odd},\\ \mathrm{C}_{\kappa\mathrm{m}}\xrightarrow{+\sigma_{z}}\mathrm{C}_{\kappa\mathrm{m}},&\text{if }\kappa\text{ is even}.\end{cases}$ (8)
Rotational and horizontal mirror symmetries constrain the composition of an eigenstate but not its absolute orientation in $\varphi$, which is fixed by vertical mirror symmetry. From Eq. Möbius-topological auxiliary function for $f$ electrons, the phase differences $\mathrm{Arg}(A_{j}A_{k}^{*})$ determine the azimuthal positions of the density maxima. For a two-component state $\left|\psi\right\rangle=C_{1}|M_{1}\rangle+C_{2}|M_{2}\rangle$, vertical mirror symmetry at $\varphi=\varphi_{\mathrm{m}}$ imposes:
@@
where the tangent removes the ambiguity of a global phase. While the extension to multi-component states is more involved, this relation illustrates how auxiliary functions accommodate vertical mirror symmetry.
@@
Example.— We consider the optically active material NdCl3 [34] as an example. The Nd3+ ion ($J=9/2$) experiences a crystal field generated by the nine surrounding Cl- ions. First, this environment exhibits a threefold rotational symmetry, yielding $\gcd(M_{j}-M_{k})=3$. Second, the presence of a horizontal mirror plane leads to $\mathrm{I}+\mathrm{C}_{3}=\mathrm{S}_{6}\xrightarrow{+\sigma_{z}}\mathrm{C}_{6{\rm m}}$, so that the auxiliary function acquires sixfold rotational symmetry with $\gcd(M_{j}-M_{k})=6$. As a result, the eigenstates take the forms:
@@
Third, taking into account the vertical mirror plane at $\varphi_{\rm m}=-6.72^{\circ}$ (see SM Fig. S1), Eq. 9 gives:
@@ Acknowledgements.
In this work, the auxiliary function $\psi(r,\theta,\varphi)$ is normalized, as are the spherical harmonics $Y_{J}^{M}(\theta,\varphi)$. Consequently, the radial function $R(r)$ in Eq. Möbius-topological auxiliary function for $f$ electrons is also normalized, i.e.,
$\int_{0}^{\infty}|R(r)|^{2}r^{2}\,\mathrm{d}r=1.$
For the Hamiltonian in Eq. (1), the radial function takes the form:
@@
where
$R_{n}^{l}(x)=\sqrt{\frac{(n-l-1)!}{2n(n+l)!}}\,x^{l}e^{-x/2}L_{n-l-1}^{2l+1}(x),$
and $L_{n-l-1}^{2l+1}(x)$ denotes the associated Laguerre polynomial. In this work, the atomic number $Z$ in the above expressions is replaced by an effective nuclear charge $\tilde{Z}$, defined as follows:
$\tilde{Z}=\frac{n}{n^{*}}Z^{*},$
@@
These four properties hold for associated Legendre polynomials with integer degrees and orders. For half-integer degrees, constructing the corresponding functions requires additional care. Properties (S2S5) correspond one-to-one to those listed in Table 1, with the second and third properties in the table weaker than those in Eqs. S3 and S4 to accommodate Hunter’s alternative normalization convention.
Eqs. S2S5 each have specific implications. Eq. S2 ensures that $Y_{J}^{M}(\theta,\varphi)$ is an eigenfunction of $\hat{L}^{2}$ with eigenvalue $J(J+1)\hbar^{2}$ and of $\hat{L}_{z}$ with eigenvalue $M\hbar$, namely:
@@
Hobson’s Legendre functions, obtained by solving the associated Legendre equation with complex parameters, come in two types: $P_{\lambda}^{\mu}(z)$ and $Q_{\lambda}^{\mu}(z)$. Among these, $P_{\lambda}^{\mu}(z)$ serves as the natural extension of the Legendre polynomial, with its explicit form given by [20, 21, 22, 23, 24, 25, 26, 27, 28]:
@@
where $\lambda$, $\mu$, and $z$ may be complex, and $_{2}F_{1}$ denotes the hypergeometric function. When $\lambda$ and $\mu$ are integers or half-integers $J$ and $M$, $P_{J}^{M}(x)$ can be expressed as follows:
@@
These results agree with those given in Eqs. S6 and S7. Table S1 presents specific forms of Hobson’s $P_{J}^{M}(x)$.
@@
Hobson’s Legendre functions satisfy Eqs. S2 and S5, but for half-integer degrees and orders they fail to satisfy Eq. S4. For half-integer orders $M\leq 1$, Eq. S3 remains valid; for $M>1$, however, it breaks down and the integral diverges. Consequently, no normalization coefficient can be defined, and the functions are intrinsically non-normalizable.
Because they fail to satisfy Eqs. S3 and S4, Hobson’s Legendre functions are unsuitable for the present work.
@@
[…]
@@
Table S2: Hunter’s $P_{J}^{M}(x)$: In Hunter’s convention, $P_{J}^{M}(x)$ may be taken as $P_{J}^{|M|}(x)$ for $M<0$.
@@
Existing Legendre functions do not satisfy all four conditions, Eqs. S2S5. For half-integer $J$ and $M$, Eqs. S4 and S5 conflict; for example, when $M=\pm 1/2$, one requires opposite parity while the other requires the same. This is resolved by introducing the factor $\operatorname{sgn}(x)$.
@@
Table S3: $P_{J}^{M}(x)$ in this work for $x=\pm|x|$, with $x\neq 0$.
@@
As stated in the text, taking NdCl3 as an example. The total angular momentum of Nd3+ is $J=9/2$. To analyze its symmetry, we first consider rotational symmetry. The nine nearest-neighbor Cl- ions surrounding Nd3+ generate a crystal field with threefold rotational symmetry. Substituting $J=9/2$ and $\kappa=3$ into Eq. 7, we obtain:
@@
Second, we examine horizontal mirror symmetry to further characterize the system. The nine nearest-neighbor Cl- ions possess horizontal mirror symmetry. From Eq. 8, with $\kappa=3$ being odd, the symmetry transforms as
$\mathrm{I}+\mathrm{C}_{3}=\mathrm{S}_{6}\xrightarrow{+\sigma_{z}}\mathrm{C}_{6\mathrm{m}},$
implying that the auxiliary function exhibits six-fold rotational symmetry. Consequently, the true eigenstates of NdCl3 form a subset of Eq. S17, adopting sixfold symmetry as follows:
[…]
For $|\pm\tfrac{1}{2}\rangle$, the difference in $M$ with any other state is less than 6. To realize sixfold rotational symmetry, $|\pm\tfrac{1}{2}\rangle$ must therefore remain isolated as a distinct state.
Third, we analyze the vertical mirror symmetry. As shown in Fig. S1, one of the vertical mirror planes of NdCl3 is located at $\varphi_{\rm m}=-6.72^{\circ}$. Using Eqs. 9 and S18, we obtain:
Figure S1: Ionic structure around Yb3+ in NdCl3, showing one of the vertical mirrors at $\varphi_{\rm m}=-6.72^{\circ}$.
Figure S2: Auxiliary densities of the eigenstates of NdCl3. They all satisfy the symmetries of the crystal field. The dashed line in each relevant subplot indicates one of the vertical mirrors, consistent with Fig. S1. $\bigl||\psi_{4}\bigr>_{\pm}\big|^{2}$ has vertical mirror symmetry at any angle. The surface represents the equiprobability density surface that encloses 50% of the total auxiliary density.
Using the auxiliary function, the results of Eqs. S18 and S19 can be derived through straightforward analysis. These results are further validated by McPhase, a software package for computing static and dynamic magnetic properties of rare-earth compounds, including phase diagrams and excitations [35]. With McPhase, the precise ground and excited states of NdCl3 can be obtained as follows:
[…]
It can be seen that all the eigenstates conform to Eq. S18, which verifies the validity of the result obtained by the auxiliary function.
@@
A comparison between Eqs. S19 and S21 shows that the prediction in Eq. S19 is highly accurate, exhibiting only a very small discrepancy from Eq. S21, likely due to the approximations in the McPhase computation.
We have plotted the auxiliary density distribution corresponding to the eigenstates of NdCl3, as shown in Fig. S2. It is evident that the distribution perfectly conforms to the symmetry inherent in the crystal field.

### v3

@@ Abstract
where $Y_{l}^{m}(\theta,\varphi)$ are spherical harmonics constructed from the associated Legendre polynomials $P_{l}^{m}(\cos\theta)$. Here, $l$ and $m$ denote the angular and magnetic quantum numbers, which also correspond to the degree and order of $P_{l}^{m}$, respectively, associated with the orbital angular momentum and its $z$-component. Throughout this paper, unless otherwise stated, $l\in\mathbb{N}$ and $m\in\mathbb{Z}$ satisfy $l\leq m\leq l$.
@@
Figure 2: (a) The factor $e^{i\varphi/2}$ is plotted along a circular arc of radius 3, with its vector trajectory forming a Möbius strip. (b) Auxiliary densities distribution of $f$-electron states in Nd3+ ($J=9/2$). Kets $|M\rangle$ denote the quantum states. The auxiliary densities exhibit D∞h symmetry, and the surface encloses 75% of the total density.
Symmetry analysis.— Auxiliary functions can be regarded as approximate wave functions, and one of their important characteristics is that the square modulus (auxiliary density) satisfies the crystal field symmetry. Thus, auxiliary functions can be employed to determine energy eigenstates, provided their densities conform to the crystal-field symmetry.
Rotational symmetry is essential in constraining the eigenstates. The auxiliary density of a pure eigenstate with a single $M$ exhibits D∞h symmetry. A superposition, however, generally breaks this symmetry, reducing it to discrete rotational symmetries.
For instance, for a two-state superposition $\left|\psi\right\rangle=C_{1}\left|M_{1}\right\rangle+C_{2}\left|M_{2}\right\rangle$, the auxiliary density is:
@@
where $A$ and $B$ are independent of $\varphi$. The auxiliary density thus has $|M_{1}-M_{2}|$-fold rotational symmetry. For example, $|M_{1}-M_{2}|=3,\,2,\,1$ give threefold, twofold, and onefold (C1) symmetry, respectively.
@@
Therefore, the auxiliary density exhibits $n_{\mathrm{rot}}$-fold rotational symmetry if and only if the following condition holds:
$n_{\mathrm{rot}}\mid\Delta M,$ (6)
where $\Delta M=M_{j}-M_{k}$ for any $j\neq k$. From Eq. 6, the eigenstate whose auxiliary density $|\psi|^{2}$ exhibits $n_{\mathrm{rot}}$-fold rotational symmetry can be written as
$|\psi\rangle=\sum_{j=0}^{\lfloor(2J-k)/n_{\mathrm{rot}}\rfloor}C_{j}\,|J-n_{\mathrm{rot}}j-k\rangle,$ (7)
where $k\in\{0,1,\dots,n_{\mathrm{rot}}-1\}$.
When analyzing rotational symmetry, it should be kept in mind that the rotational symmetry order of the auxiliary density does not necessarily coincide with that of the crystal field; the former can be higher than the latter. This originates from the fact that the auxiliary density is always inversion-symmetric (see proof in SM). In particular, when the crystal field has an odd-fold rotational symmetry and a mirror plane perpendicular to the rotation axis, the rotational symmetry order of the auxiliary density becomes twice that of the crystal field. As an example, for a crystal field with $\mathrm{C}_{n\mathrm{h}}$ symmetry, the symmetry of the auxiliary density falls into the following two cases:
$\mathrm{C}_{\mathrm{i}}\times\mathrm{C}_{n\mathrm{h}}=\begin{cases}\mathrm{C}_{2n\mathrm{h}},&\text{if }n\text{ is odd},\\ \mathrm{C}_{n\mathrm{h}},&\text{if }n\text{ is even}.\end{cases}$ (8)
Strictly speaking, $n_{\mathrm{rot}}$ in Eq. 7 denotes the rotational symmetry order of the auxiliary density.
Beyond the analysis above, additional constraints can be obtained by exploiting other symmetries of the crystal field. Symmetry considerations can substantially restrict the allowed form of the eigenstates, in some cases even fixing them uniquely, depending on the specific situation. Further analysis may proceed by incorporating vertical mirror symmetries, which impose constraints on the relative phases of the coefficients appearing in the eigenstate expansion.
From Eq. Möbius-topological auxiliary function for $f$ electrons, the phase differences $\mathrm{Arg}(A_{j}A_{k}^{*})$ determine the azimuthal positions of the density maxima. For a two-component state $\lvert\psi\rangle=C_{1}\lvert M_{1}\rangle+C_{2}\lvert M_{2}\rangle$, a vertical mirror plane located at $\varphi=\varphi_{\mathrm{m}}$ imposes the condition
@@
where the use of the tangent removes the ambiguity associated with an overall global phase. Although the generalization to multi-component states is more involved, this example illustrates how auxiliary functions encode vertical mirror symmetry in a transparent manner.
@@
Example.— We consider the optically active material NdCl3 [34] as an example. The Nd3+ ion ($J=9/2$) is subject to a crystal field produced by nine surrounding Cl- ions, forming a local environment of $\mathrm{D}_{3\mathrm{h}}$ symmetry. Because the auxiliary density is always inversion-symmetric, the point group $\mathrm{D}_{3\mathrm{h}}$, upon combination with $\mathrm{C}_{\mathrm{i}}$ in a manner analogous to Eq. 8, is elevated to $\mathrm{D}_{6\mathrm{h}}$. Consequently, the rotational symmetry order is $n_{\mathrm{rot}}=6$. From Eq. 6, the eigenstates take the forms:
@@
Taking into account the vertical mirror plane located at $\varphi_{\mathrm{m}}=-6.72^{\circ}$ (see SM Fig. S1), Eq. 9 yields:
@@ Acknowledgements.
In this work, the auxiliary function $\psi(r,\theta,\varphi)$ and the spherical harmonics $Y_{J}^{M}(\theta,\varphi)$ are both normalized. Consequently, the radial function $R(r)$ in Eq. Möbius-topological auxiliary function for $f$ electrons is also normalized, i.e.,
$\int_{0}^{\infty}|R(r)|^{2}r^{2}\,\mathrm{d}r=1.$ (S1)
For the Hamiltonian in Eq. Möbius-topological auxiliary function for $f$ electrons, the radial function takes the form
@@
where $n$ denotes the principal quantum number, $Z$ is the nuclear charge, and $a_{0}$ is the Bohr radius. The dimensionless radial function is given by
$R_{n}^{l}(x)=\sqrt{\frac{(n-l-1)!}{2n(n+l)!}}\,x^{l}e^{-x/2}L_{n-l-1}^{2l+1}(x),$ (S3)
where $L_{n-l-1}^{2l+1}(x)$ denotes the associated Laguerre polynomial.
In this work, $Z$ is replaced by an effective nuclear charge $\tilde{Z}$, defined as
$\tilde{Z}=\frac{n}{n^{*}}Z^{*},$ (S4)
@@
These four properties hold for associated Legendre polynomials with integer degrees and orders. For half-integer degrees, constructing the corresponding functions requires additional care. Properties (S5S8) correspond one-to-one to those listed in Table 1, with the second and third properties in the table weaker than those in Eqs. S6 and S7 to accommodate Hunter’s alternative normalization convention.
Eqs. S5S8 each have specific implications. Eq. S5 ensures that $Y_{J}^{M}(\theta,\varphi)$ is an eigenfunction of $\hat{L}^{2}$ with eigenvalue $J(J+1)\hbar^{2}$ and of $\hat{L}_{z}$ with eigenvalue $M\hbar$, namely:
@@
Hobson’s Legendre functions arise from the associated Legendre equation with complex parameters and come in two types: $P_{\lambda}^{\mu}(z)$ and $Q_{\lambda}^{\mu}(z)$. Among these, $P_{\lambda}^{\mu}(z)$ serves as the natural extension of the Legendre polynomial, with its explicit form given by [20, 21, 22, 23, 24, 25, 26, 27, 28]:
@@
where $\lambda$, $\mu$, and $z$ may be complex, and $_{2}F_{1}$ denotes the hypergeometric function. When $\lambda$ and $\mu$ take integer or half-integer values $J$ and $M$, $P_{J}^{M}(x)$ can be expressed as follows:
@@
These expressions are consistent with Eqs. S9 and S10 where applicable. Table S1 lists explicit forms of Hobson’s $P_{J}^{M}(x)$.
@@
Hobson’s Legendre functions satisfy Eqs. S5 and S8, but for half-integer degrees and orders they fail to satisfy Eq. S7. For half-integer orders $M\leq 1$, Eq. S6 remains valid; for $M>1$, however, it breaks down and the integral diverges. Consequently, no normalization coefficient can be defined, and the functions are intrinsically non-normalizable.
Because they fail to satisfy Eqs. S6 and S7, Hobson’s Legendre functions are unsuitable for the present work.
@@
Hunter gave $P_{J}^{M}$ only for $M\geq 0$, as negative $M$ is redundant under his convention $Y_{J}^{M}\equiv P_{J}^{|M|}(\cos\theta)e^{iM\varphi}$[30], so one can define $P_{J}^{-|M|}\equiv P_{J}^{|M|}$. This ensures the third condition in Table 1, and the integrability of $|P_{J}^{M}|^{2}$ satisfies the second condition. Although Hunter’s functions do not exactly match Eqs. S6 and S7 due to coefficient conventions, their integrability and symmetry remain valid.
@@
Table S2: Hunter’s $P_{J}^{M}(x)$. $P_{J}^{M}(x)$ can be taken as $P_{J}^{|M|}(x)$ for $M<0$.
@@
Existing Legendre functions do not satisfy all four conditions, Eqs. S5S8. For half-integer $J$ and $M$, Eqs. S7 and S8 conflict; for example, when $M=\pm 1/2$, one requires opposite parity while the other requires the same. This is resolved by introducing the factor $\operatorname{sgn}(x)$.
@@
Table S3: $P_{J}^{M}(x)$ in this work for $x=\pm|x|$ $(x\neq 0)$.
@@
We provide additional details underlying the example discussed in the main text.
Starting from the inversion symmetry of the auxiliary density, the combination $\mathrm{C}_{\mathrm{i}}\times\mathrm{D}_{3\mathrm{h}}=\mathrm{D}_{6\mathrm{h}}$ implies a sixfold rotational symmetry ($n_{\mathrm{rot}}=6$). Substituting $J=9/2$ into Eq. 7 yields the symmetry-allowed eigenstate structures:
$\lvert\psi\rangle=\begin{cases}C_{\pm\frac{9}{2}}\lvert\pm\frac{9}{2}\rangle+C_{\mp\frac{3}{2}}\lvert\mp\frac{3}{2}\rangle,\\[2.0pt] C_{\pm\frac{7}{2}}\lvert\pm\frac{7}{2}\rangle+C_{\mp\frac{5}{2}}\lvert\mp\frac{5}{2}\rangle,\\[2.0pt] \lvert\pm\frac{1}{2}\rangle.\end{cases}$ (S20)
The states $\lvert\pm\tfrac{1}{2}\rangle$ remain isolated, as their magnetic quantum numbers differ by less than six from all other basis states.
Imposing the vertical mirror symmetry at $\varphi_{\mathrm{m}}=-6.72^{\circ}$ (Fig. S1) and using Eq. 9, one obtains
These results are further validated by McPhase, a software package for computing static and dynamic magnetic properties of rare-earth compounds [35]. With McPhase, the precise ground and excited states of NdCl3 can be obtained as follows:
@@
All eigenstates conform to Eq. S20, confirming the validity of the auxiliary-function result.
@@
A comparison between Eqs. S21 and S23 shows that the prediction in Eq. S21 is highly accurate, exhibiting only a very small discrepancy from Eq. S23, likely due to the approximations in the McPhase computation.
We have plotted the auxiliary density distribution corresponding to the eigenstates of NdCl3, as shown in Fig. S2. The distribution fully respects the symmetry of the crystal field.
Figure S1: Ionic structure around Nd3+ in NdCl3, showing one of the vertical mirrors at $\varphi_{\rm m}=-6.72^{\circ}$.
Figure S2: Auxiliary densities of the eigenstates of NdCl3. They all satisfy the symmetries of the crystal field. The dashed line in each relevant subplot indicates one of the vertical mirrors, consistent with Fig. S1. $\bigl||\psi_{4}\bigr>_{\pm}\big|^{2}$ has vertical mirror symmetry at any angle. The surface represents the equiprobability density surface that encloses 75% of the total auxiliary density.